5. Use trigonometry to find all sides and angles in the triangles below 10 in 6 m 35 3 m
Answer #1

10 in 6 m 35 3 m

a) In triangle ABC, Given that AC 10 in , B-90 , С 350

sin 350-AB AC AB AC sin 35 , that is AB 10 sin 35° 5.74 in

BC AC cos 35° cos35° = AC cos 350 = , that is BC-10 cos 35°-8. 19 in

angle A+angle B+angle C=180degree

LA + 90° 350-180.

LA 55°

----------------

b) In triangle PQR , OR = 3 m , PR=6,m , R- 90

From Pythagoras theorem , PQ^2=QR^2+PR^2

PQ^2=3^2+6^2Rightarrow PQ=6.71,m

sinangle Q=rac{PR}{PQ}Rightarrow angle Q=sin^{-1}left ( rac{PR}{PQ} ight )=sin^{-1}left ( rac{6}{6.71} ight )=63.4degree

angle P+angle Q+angle R=180degree

angle P+63.4degree+90degree=180degree

Rightarrow angle P=26.6degree

-------------------